OpenStudy (anonymous):

Which is equivalent to (7 – 3)4?

OpenStudy (anonymous):

(7 – 3)4=7*4-3*4=28-12=16

Directrix (directrix):

@liltony987 Is this --> (7 – 3)4 supposed to be (7 – 3)^4 ? Just wondering.